OpenStudy (anonymous):

Simplify (2√5+3√7)^2. Justify your answer and Show Your Work! :) Thanks!

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (a + b)^2 = a^2 + b^2 + 2ab }\) Use this identity to solve.

OpenStudy (anonymous):

So... The first step of this equation should be writing it like this; (2√5+3√7)^2=2√5^2+3√7^2+2(2√5 3√7) Sorry if I'm confusing anyone, I'm horrible at this stuff...

OpenStudy (anonymous):

Well, the first step is (2√5+3√7)^2=(2√5)^2+(3√7)^2+2(2√5 3√7) then after simplifying (2√5+3√7)^2= 4 * 5 + 9 * 7 + 12 * √35 = 20 + 63 +12 √35 = 83 + 12√35